27% of 350= 12% of 8= 7% of 350=
Answer by tutorcecilia(2152)

Change the percent to a decimal by dividing it by 100 and multiply:
.
27% of 350= 27/100 X 350 = .27 X 350 = 94.5
.
12% of 8 = 12/100 X 8 = .12 X 8 = .96
.
7% of 350 = 7/100 X 350 = .07 X 350 = 24.5